1. Calculate the density (in g/L) of Freon-12, CF2Cl2, at 27.5°C and 0.949 atm.
2. A mixture of 0.222 g of H2, 1.17 g of N2, and 0.828 g of Ar is stored in a closed container at STP. Find the volume (in L) of the container, assuming that the gases exhibit ideal behavior.
3. What volume (in L) of O2 at STP is required to oxidize 9.2 L of NO at STP to NO2?
WebAssign will check your answer for the correct number of significant figures.(Response) L
What volume (in L) of NO2 is produced at STP?
WebAssign will check your answer for the correct number of significant figures.(Response) L
Answer #1

Sol 1.

As P = \rho R T / M

where P = Pressure = 0.949 atm

\rho = Density  

R = Gas constant = 0.0821 L atm / K mol  

T = Temperature = 27.5 °C = 27.5 + 273 = 300.5 K

M = molar mass of CF2Cl2 = 120.91 g/mol

So , \rho = PM / RT  

= ( 0.949 × 120.91 ) / ( 0.0821 × 300.5 )

=   4.651 g/L